# Entanglement engineering of optomechanical systems by reinforcement learning

### v1

@@ IV Entanglement engineering in nonlinear quantum optomechanics
Figure 7: Generating target for deep-RL based creation and stabilization of entanglement in a nonlinear open quantum optomechanical system. The nonlinear interaction strength is $g_{0}=0.2\,\omega_{m}$. The PPO agent creates entanglement characterized by $E_{N}\sim\log 2$ versus time, calculated through the SME with dissipation to the vacuum bath for $\kappa=0.1\,\omega_{m}$, and $\gamma=0.01\,\kappa$. The system is initialized in the vacuum state $|\psi\rangle=|00\rangle$, i.e., the pure state, with $10\times 10$ Fock bases. There are then ten stochastic Fock projectors $\hat{P}_{n}=|n\rangle\langle n|$ with the respective Fock numbers $n=0,1,\ldots,9$, in the SME with $\eta=0.1$. The time-dependent control signal is the detuning $\Delta$ and the amplitude $\alpha_{L}$ of the driven laser within the fixed range $\Delta,\alpha_{L}\in[-5,5]\,\omega_{m}$. (a,b) Trained quantities $\widetilde{R}$ and $\widetilde{E}_{N}$ converge to a certain value as the episode number increases, as illustrated by the light-color curves, where the dark blue and orange traces represent the data averaged over 100 previously consecutive episodes. (c,d) Time-dependent series of $E_{N}(t)$ and the driven laser signals $\Delta,\alpha_{L}$ at a certain episode selected from the training converged regime in (a,b). (e,f) The corresponding photon and phonon statistics on the Fock basis at the end of the time point of the selected training episode in (c,d). (g) The time evolution of the corresponding expected measurement current, including the expected number $\langle n_{p}\rangle$ and phonon $\langle n_{m}\rangle$ of photons in the Fock basis, where the time series of $\langle n_{p}\rangle$(t) serves as the target to construct the reward function in the next phase, i.e., the experimental version shown in Fig. 8. At each step, the observation is $E_{N}(t)$. The reward function is constructed as $R(t)=-|E_{N}(t)-\log 2|-|\langle n_{p}\rangle(t)+\langle n_{m}\rangle(t)-a|/b$, where the purpose of the second term is to limit the range of excited quantum states in the Fock basis by the total number $\langle n_{p}\rangle+\langle n_{m}\rangle$ with numerically optimized hyperparameters $a=1$ and $b=30$.
@@
Figure 8: Entanglement engineering by the recurrent PPO agent. The target generated as described in Fig. 7 is exploited to create entanglement by $E_{N}\sim\log 2$ from the only partial observation of the expected photon number $\langle n_{p}\rangle(t)$. The reward function is $R(t)=-|\langle n_{p}\rangle(t)-\langle n_{p}^{\textnormal{target}}\rangle(t)|$, where the target time series of $\langle n_{p}^{\textnormal{target}}\rangle(t)$ is from the target-generating process in Fig. 7(g). In this training configuration, while only partial information is extracted from the system, the performance measures in (a-g) display a similar behavior compared with those in Fig. 7. Other aspects of the setting and parameters are the same as in Fig. 7.
@@
In an open quantum optomechanical system under the strong laser-driven approximation, the radiation pressure on the movable mechanical mirror generates a linear interaction between the optical and mechanical modes. When this approximation does not hold, the interaction between the two modes becomes nonlinear. Entanglement can still be created despite the nonlinear interaction, but control becomes more challenging. In particular, in the standard quantum optomechanical system described by Eq. (II.1), the nonlinear coupling term $\hbar g_{0}\hat{a}^{\dagger}\hat{a}(\hat{b}^{\dagger}+\hat{b})$ can be used to create entanglement, but high-level quantum states can also be excited during the process, making it difficult to stabilize the entanglement within a finite Fock basis. Realistically, the quantum dynamics are governed by the SME due to the WCM, which induces the nonlinear stochastic evolution. The problem then becomes that of creating and stabilizing the entanglement of non-Gaussian states decaying to the vacuum bath. Despite the difficulties, deep-RL can still provide a general approach through some optimal combination of the neural network structure, observable, reward function, and action.
@@
We consider the nonlinear optomechanical system described by Eq. (II.1) and exploit deep RL to set the control goal of achieving the entanglement near $E_{N}\sim\log 2$. This desired entangled state is similar to the maximum entangled Bell state in the corresponding linear system. For entanglement engineering of a nonlinear optomechanical system, a key issue is how to choose an effective and experimentally feasible observation quantity. With the general actor and critic neural network, the deep-RL agent is able to learn how to harness the entanglement of the optomechanical system by utilizing a laser with only partial state observation. More specifically, to achieve control, we articulate a training process consisting of two phases: target-generating and target-utilization phases, as schematically illustrated in Fig. 6.
@@
The first training step is the target-generating phase, in which numerical simulations are used to generate the observation and reward data and the PPO agent observes the logarithmic negativity $E_{N}(t)$ and constructs the reward function combining the expectation number of the photons and phonons:
$R(t)=-|E_{N}(t)-\log 2|-|\langle n_{p}\rangle(t)+\langle n_{m}\rangle(t)-a|/b.$ (26)
Figure 7 shows the control results, where the excitation of quantum states is limited by the total number $\langle n_{p}\rangle+\langle n_{m}\rangle$. The target time series of the expected photon number is obtained as $\langle n_{p}^{\textnormal{target}}\rangle(t)$. The second step is the target-utilization phase, during which the reward function is the target $R(t)=-|\langle n_{p}\rangle(t)-\langle n_{p}^{\textnormal{target}}\rangle(t)|$. Since it is time-dependent, the recurrent neural network added after the MLPs displays a strong and stable learning ability, which outperforms the case with only MLPs. The expected photon number $\langle n_{p}\rangle(t)$ is observed by the recurrent PPO agent as
$\langle n_{p}\rangle=\sum_{n}n\langle\hat{P}_{n}\rangle,$ (27)
@@
which is more experimentally feasible than the quantity $E_{N}$. While the recurrent neural network has some considerable advantages as discussed in Sec. D.2, it still encounters the challenge of engineering optimization [40] in order to achieve a correct and efficient implementation. In our case, the main challenge is the time cost to optimize the parameters to search for a global minimum or maximum due to the ten stochastic collapse operators, $\hat{P}_{n}$ with $n=0,1,\ldots,9$, requiring a long simulation time. Our solution is to consider only the $\mathbb{N}=1$ quantum optomechanical environment, in which the agent collects data and updates the policy every $\mathbb{Z}=15$ and $\mathbb{Z}=5$ episodes in two phases (target-generating and target-utilization), respectively, with the time horizon $T=500$. Note that, using ten stochastic projectors $\hat{P}_{n}$ can result in a large variance in the WCM flow:
$\mathcal{I}^{\hat{P}_{n}}(t)=\sum_{n}n\left[\langle\sqrt{\eta}\hat{P}_{n}\rangle+\frac{dW_{n}(t)}{\sqrt{4\eta}dt}\right]$ (28)
where ten independent Wiener processes $dW_{n}(t)$ are used. In this case, observation of the measured random current is infeasible.
@@
Representative results are as follows. In the target-generating phase, despite the disturbance of the stochastic process from WCM, the training curves for both the reward $\widetilde{R}$ and the logarithmic negativity $\widetilde{E}_{N}$ converge with the episode number, as shown in Figs. 7(a,b), indicating that entanglement has been created and stabilized by the well-trained PPO agent, as shown by Fig. 7(c) with the laser control signal displayed in Fig. 7(d). At the end of the time period, the photon and phonon statistics with respect to the Fock basis are shown in Figs. 7(e,f), where the reduced photon state exhibits an oscillating tail that resembles the displaced squeezed state and the reduced phonon state displays the thermal-like state. Figure 7(g) shows the corresponding target pattern $\langle n^{\textnormal{target}}_{p}\rangle(t)$. In the target-utilization phase, the recurrent PPO agent is able to steadily learn to create and stabilize the entanglement, as shown in Fig. 8, where only partial information is extracted from the quantum optomechanical environment. Especially, various entangled states have been created, such as a reduced photon state with the head oscillating on the Fock basis in photon statistics entangled with the thermal-like reduced phonon state, as exemplified in Figs. 8(e,f). Due to the nonlinear and stochastic process in the SME, the entangled states created and controlled are non-steady states, rendering infeasible Bayesian control. We thus employ random control as a benchmark, where a uniformly random distribution of actions is taken in a certain range $\Delta,\alpha_{L}\in[-5,5]\omega_{m}$ and the tested values of the measurement rate are $\eta=[0.05,0.1,0.3,0.5,0.7,1]$. Figure 9 shows that, as the measurement rate increases, the random control is unable to harness the entanglement while our well-trained recurrent PPO agent can maintain the entanglement percentage at $50\%$ or higher.
@@
V Physical understanding of reinforcement-learning based entanglement engineering
@@
V.1 Linear interaction regime
For the linear system described by Eq. (3), the maximum entanglement corresponds to a Bell state, of which the expected photon number is 0.5. Intrinsically, the beam splitter Hamiltonian is capable of generating Bell states [85, 86, 87], a reasonable assumption is that, when the expected photon number reaches the value of 0.5, the maximum entanglement is achieved in a linearly interacting quantum optomechanical system. This assumption provides the base for constructing the reward function $R(t)=-|\langle\hat{n}_{p}\rangle(t)-0.5|$, where the deviation in the expected photon number from 0.5 results in a decreasing reward and therefore implies reduced entanglement. As illustrated in Fig. 10, the RL agent is designed to maximize the accumulated reward value, which is equivalent to stabilizing the expected photon number about the value of 0.5 for as long as possible. The testing results shown in Fig. 3 indicate that the maximum entanglement can indeed be created and stabilized by the RL control.
@@ V.1 Linear interaction regime
A central step in RL control is to modulate the laser input based on the measured photon number, which requires the relationship between the laser driving and the photon number. When the frequency of the laser is in the red-detuned regime: $\Delta=\omega_{L}-\omega_{c}=-\omega_{m}$, the quantum state switches between the two modes - the cavity optical and the mechanical oscillator modes, leading to a “swap” Hamiltonian. The coefficient $G$ in Eq. (3) is proportional to the amplitude of the cavity parameter $\bar{\alpha}_{c}$ that is determined by the laser. In the linear interaction regime, RL control is achieved via two adjustments of the laser based on the measured photon number: (1) the laser frequency is changed into the red-detuned regime and (2) the laser amplitude is perturbed to modulate the driving strength $G$ to control the two modes of switching, which affects the expected photon number. Note that, during this process, there is no energy gain: there is energy loss due to the dissipation of the cavity and oscillator modes into the vacuum bath, as described by Eq. (II.2) with the dissipation rate given by $\gamma=10^{-2}\kappa$. This relation means that the energy loss due to the oscillator mode occurs more slowly than that with the cavity mode. In essence, the working of the laser is to transfer the energy from the oscillator mode to the cavity mode to stabilize the photon number to a desired value. The underlying dissipation process is not beneficial to the entanglement, as it cannot be modulated by the “swap” term in the Hamiltonian, eliminating any possibility of entanglement enhancement in an optomechanical system in the linear interaction regime. It is worth noting that, in the nonlinear interaction regime, entanglement enhancement is possible, as will be described below.
V.2 Nonlinear interaction regime
@@
When the interactions between the optical and mechanical modes are nonlinear, the relationship between entanglement and photon number can be sophisticated and is currently unknown. However, RL can be used to find the relation numerically. To achieve this, we first assume that there is a solution of the one-to-one correspondence between $E_{N}$ and $\langle\hat{n}_{p}\rangle$ in the time domain. The reward function is constructed according to $E_{N}$ to train the RL agent to maximize the accumulated reward. In the testing phase, the time evolution of the expected photon number is obtained, leading to the target time series of the expected photon number, as exemplified in Fig. 7. Note that the “best” photon number is no longer simply 0.5: it is now time-dependent. In the next episode, the reset RL agent will learn to control the system based on the target’s expected photon number. The performance of the new RL agent in the testing phase, as shown in Fig. 8, validates our initial assumption about the existence of the one-to-one correspondence between $E_{N}$ and $\langle\hat{n}_{p}\rangle$, even though it is time-dependent.
@@
In the nonlinear interaction regime, the physical picture of how the laser leverages the radiation-pressure interaction to create and stabilize the photon number and even the entanglement is not straightforward. However, physical insights can be gained by examining the strong laser limit. When the amplitude of the laser is strong: $\bar{\alpha}_{c}\gg 1$, in the blue-detuned regime with $\Delta=+\omega_{m}$, the laser can modulate its frequency to create exponential growth of the energies of both the cavity and oscillator modes, accompanied by the generation of strong quantum correlation between the two modes. In the red-detuned regime with $\Delta=-\omega_{m}$, a switching process between the two modes occurs, which is the same as that in the linear interaction regime described in Sec. V.1.
@@
The blue- and red-detuned regimes have a competitive relationship with each other in terms of both the photon number and entanglement. In particular, in the blue-detuned regime, photons are excited and the rate of excitation can be larger than that associated with quantum dissipation to the vacuum bath. In addition, quantum entanglement is enhanced, overcoming decoherence from the classical environment. However, in the red-detuned regime, no photons are excited and there is only a two-mode energy-transferring process that does not completely suppress the process of quantum dissipation to the vacuum bath, resulting in photon loss and eventually reducing entanglement. Stabilizing the photon number and entanglement requires a balance between the operations in the blue- and red-detuned regimes. In general, the blue-detuned regime is prone to too high photon levels with strong entanglement, which should be balanced by the red-detuned regime operation to reduce the photon number, as shown schematically in Fig. 11. Overall, in the nonlinear interaction regime, laser driving of finite amplitude and frequency modulation can control the photon number and entanglement to a certain extent. An example is shown in Fig. 8(d), where the RL agent finds the optimal action flow with a finite laser amplitude. Note that, the detuning $\Delta$ is modulated mainly in the range $\Delta\in[-2\omega_{m},2\omega_{m}]$, signifying a balance between the blue- and red-detuned operations.
V.3 Weak continuous measurement
@@
In an open quantum system, under WCM and quantum dissipation into the vacuum bath as well, a Wiener process occurs in the observable. More specifically, the Wiener process arises from the Gaussian-weighted projection over the eigenstates [Eq. (49) in Appendix B], which weakly extracts the partial information from the quantum system and induces stochastic disturbances in both the dynamical equation [Eq. (II.2)] and observation [Eq. (10)]. Such disturbances have the advantage of avoiding a complete quantum state collapse and providing the capability to extract the quantum information continuously in the time domain. However, the nonlinear stochastic process occurs in both quantum dynamical trajectories and the measurement current, making it challenging to continuously control the quantum system through WCM.
@@
For stochastic noise in the WCM current, the present cutting-edge technology enables the extraction of quantum information by the RL agent through a process resembling noise filtering. Specifically, the observation in the reward function Eq. (12) is the WCM current. We can use a Gaussian filter for data pre-processing. Further, $n_{traj}$ quantum ensembles are utilized to reduce the variance. In addition, the RL agent is trained to maximize the accumulated reward, which serves to average the stochastic term in the measurement current over time. These noise filtering processes are helpful for extracting the information about the expected photon number and the information about the target quantum state. For the nonlinear quantum stochastic process with quantum dissipation in Eq. (II.2), the RL agent successfully trains the laser to leverage the interactions between the optical and mechanical modes, linear or nonlinear, to mitigate quantum decoherence and dissipation, as exemplified in Figs. 3 and 8, respectively.
VI Discussion
Summary of main results.

### v2

@@ Results
Figure 6: Generating target for deep-RL based creation and stabilization of entanglement in a nonlinear open quantum optomechanical system. (a,b) Trained quantities $\widetilde{R}$ and $\widetilde{E}_{N}$ converge to a certain value as the episode number increases, as illustrated by the light-color curves, where the dark blue and orange traces represent the data averaged over 100 previously consecutive episodes. (c,d) Time-dependent series of $E_{N}(t)$ and the driven laser signals $\Delta,\alpha_{L}$ at a certain episode selected from the training converged regime in (a,b). (e,f) The corresponding photon and phonon statistics on the Fock basis at the end of the time point of the selected training episode in (c,d). (g) The time evolution of the corresponding expected quantities, including the expected numbers $\langle\hat{n}_{p}\rangle$ and $\langle\hat{n}_{m}\rangle$ in the Fock basis, where the time series of $\langle\hat{n}_{p}\rangle$(t) serves as the target to construct the reward function in the next phase, i.e., the experimental version shown in Fig. 7.
Experimentally, mixed quantum states are more realizable than pure states due to the quantum decoherence with the classical environment, e.g., the vacuum bath. To address this issue, and referring to the previous work [87], we assume that the initial state is a mixed state in the form of $\rho=(1-p)|10\rangle\langle 10|+p|01\rangle\langle 01|$, where the parameter $p$ is fixed or a random variable $p\in[0,1]$ because of the coupling to the classical environment. The beam-splitter Hamiltonian stipulates that the photon and phonon modes are symmetric to each other, allowing $p$ to be rescaled to the interval $p\in[0,0.5]$. Figure 5(a) shows the performance with respect to the initial mixed quantum state with the same parameter $p$ for each training and testing episode (solid traces), where the complete mixed case with $p=0.5$ leads to the worst performance but still possesses entanglement to a significant extent. The reason lies in the inherent property of the beam-splitter Hamiltonian, which can create the maximum entangled states: $[|10\rangle+e^{i\varphi}|01\rangle]/\sqrt{2}$, with respect to the part of the initial quantum state, such as $|10\rangle$ or $|01\rangle$ through the linear interactions, regardless of whether it acts on a pure or a mixed state. In Fig. 5 (a), the dashed traces display the performance during the training phase with a random initial mixed quantum state, which is generated by the random variable $p$ with the uniform distribution in the range of $p\in[0,0.5]$. The error bar characterizes the uncertainty over ten end-training episodes.
Figure 5(b) shows the testing performance of two kinds of trained models, one trained by the initial state $|\psi\rangle=|10\rangle$ and another by the random initial mixed-state $\rho=(1-p)|10\rangle\langle 10|+p|01\rangle\langle 01|$ (distinguished by dark and light colors, respectively). Note that the beam-splitter Hamiltonian transforms the initial state $|10\rangle$ or $|01\rangle$ to a Bell state with the corresponding expected photon number: $\langle\hat{n}^{\textnormal{target}}_{p}\rangle=0.5$, where the dissipative degree to the vacuum bath is much weaker than the beam-splitter interaction. However, if the initial state is the mixed state, the $|10\rangle\langle 10|$ and $|01\rangle\langle 01|$ components will become independently entangled, resulting in the total quantum state being a mixture of two entangled Bell states. As a result, a nontrivial entanglement value is expected for the initial mixed state governed by the beam-splitter Hamiltonian. With the mixed probability $p=0.5$, it results in an equal mixture of the Bell states, as shown in Fig. 5. In the testing phase, the two trained models use the same initial state for a fixed value of $p$. The two models have a comparable performance, suggesting that the deep RL method is robust to the initial randomness in a mixed state. More specifically, during the testing phase, the observation is the expected photon number or WCM photocurrent. The worse performance occurs for $p=0.5$ and for other values of $p$, the performance is symmetric about $p=0.5$ due to symmetric role of the photon and photon modes in the beam splitter Hamiltonian. Note that the model trained with the observation being the measurement photocurrent displays a small difference in the performance measure [$\langle E_{N}\rangle/E_{0}$ over the whole probability interval $p\in[0,1]$] between the best and worst cases, with less uncertainties than the case where the observation is the expected photon number. Taken together, our deep-RL model trained by the weak measurement photocurrent holds a lower mean performance but possesses robustness against mixed quantum states compared with the scenario based on observing the expected value of the photon number, due to the strong capability of RL in learning randomness and executing accurate high-dimensional data-fitting.
RL in nonlinear quantum optomechanics
@@
Figure 7: Entanglement engineering by the recurrent PPO agent. The target generated as described in Fig. 6 is exploited to create entanglement by $E_{N}\sim\log 2$ from the only partial observation of the expected photon number $\langle\hat{n}_{p}\rangle(t)$. The reward function is $R(t)=-|\langle\hat{n}_{p}\rangle(t)-\langle\hat{n}_{p}^{\textnormal{target}}\rangle(t)|$, where the target time series $\langle\hat{n}_{p}^{\textnormal{target}}\rangle(t)$ is from the target-generating process in Fig. 6(g). In this training configuration, while only partial information is extracted from the system, the performance measures in (a-g) display a similar behavior compared with those in Fig. 6. Other aspects of the setting and parameters are the same as in Fig. 6.
@@
In an open quantum optomechanical system under the strong laser-driven approximation, the radiation pressure on the movable mechanical mirror generates a linear interaction between the optical and mechanical modes. When this approximation does not hold, the interaction between the two modes becomes nonlinear. Entanglement can still be created despite the nonlinear interaction, but control becomes more challenging. In particular, in the standard quantum optomechanical system, the nonlinear coupling term $\hbar g_{0}\hat{a}^{\dagger}\hat{a}(\hat{b}^{\dagger}+\hat{b})$ can be used to create entanglement, but high-level quantum states can also be excited during the process, making it difficult to stabilize the entanglement within a finite Fock basis. Realistically, the quantum dynamics are governed by the SME due to the WCM, which induces the nonlinear stochastic evolution. The problem then becomes that of creating and stabilizing the entanglement of non-Gaussian states decaying to the vacuum bath. Despite the difficulties, model-free deep-RL can still provide a general approach through some optimal combination of the neural network structure, observable, reward function, and action.
@@
We consider the nonlinear optomechanical system and exploit deep RL to set the control goal of achieving the entanglement near $E_{N}\sim\log 2$. This nonlinear entangled state shares a similar entanglement value with the maximum entangled Bell state in the corresponding linear system. For entanglement engineering of a nonlinear optomechanical system, a key issue is selecting an effective and experimentally feasible observation quantity. Utilizing a general actor and critic neural network, the deep RL agent can learn the relationship between entanglement and the experimental observables of the optomechanical system in a model-free manner. To achieve control, we articulate a training process consisting of two phases: the target-generating phase and the target-utilization phase, facilitated by deep RL.
@@
The first training step is the target-generating phase, in which numerical SME simulations are used to generate the observation and reward data and the PPO agent interacts with the quantum environment, observes the logarithmic negativity $E_{N}(t)$ and constructs the reward function combining the expectation number of the photons and phonons: $R(t)=-|E_{N}(t)-\log 2|-|\langle\hat{n}_{p}\rangle(t)+\langle\hat{n}_{m}\rangle(t)-a|/b$ with numerically optimized hyperparameters $a=1$ and $b=30$. (Note that direct experimental measurement of the logarithmic negativity is currently not available.) Figure 6 shows the control results, where the excitation of quantum states is limited by the total number $\langle\hat{n}_{p}\rangle+\langle\hat{n}_{m}\rangle$. The target time series of the expected photon number is obtained as $\langle\hat{n}_{p}^{\textnormal{target}}\rangle(t)$. The second step is the target-utilization phase, during which the reward function is $R(t)=-|\langle\hat{n}_{p}\rangle(t)-\langle\hat{n}_{p}^{\textnormal{target}}\rangle(t)|$.
@@
Since it is time-dependent, the recurrent neural network added after the MLPs in the PPO agent displays a strong and stable learning ability, which outperforms the case with only MLPs. The expected photon number $\langle\hat{n}_{p}\rangle(t)$ is observed by the recurrent PPO agent as $\langle\hat{n}_{p}\rangle=\sum_{n}n\langle\hat{P}_{n}\rangle$, which is experimentally more feasible than the quantity $E_{N}$. While the recurrent neural network has some considerable advantages, such as long-term momery [57], it still encounters the challenge of engineering optimization [88] in order to achieve a correct and efficient implementation. In our case, the main challenge is the time cost to optimize the parameters to search for a global minimum or maximum due to the ten stochastic collapse operators, $\hat{P}_{n}=|n\rangle\langle n|$ with the respective Fock numbers $n=0,1,\ldots,9$, in the SME with the measurement rate $\eta=0.1$, requiring a long simulation time. Our solution is to consider only the $\mathbb{N}=1$ quantum optomechanical environment, in which the agent collects data and updates the policy every $\mathbb{Z}=15$ and $\mathbb{Z}=5$ episodes in two phases (target-generating and target-utilization), respectively, with the time horizon $T=500$. Note that, using ten stochastic projectors $\hat{P}_{n}$ can result in a large variance in the WCM photocurrent:
$\sqrt{\eta}\,\mathcal{I}(t)=\sum_{n}n\left[\langle\sqrt{\eta}\hat{P}_{n}\rangle+\frac{dW_{n}(t)}{\sqrt{4\eta}dt}\right],$
where ten independent Wiener processes $dW_{n}(t)$ are used. In this case, observation of the measured random photocurrent is infeasible. Even if the deep RL agent is trained in two phases with the expected photon number, it can fail during the training process due to the numerical cutoff in the Hilbert space dimension and the strong randomness introduced by the SME. In the nonlinear quantum optomechanical system, the interaction strength is $g_{0}=0.2\,\omega_{m}$. The PPO agent creates entanglement characterized by $E_{N}\sim\log 2$ versus time, calculated through the SME with dissipation to the vacuum bath for $\kappa=0.1\,\omega_{m}$, and $\gamma=0.01\,\kappa$. The system is initialized in the vacuum state $|\psi\rangle=|00\rangle$, i.e., the pure state, with $10\times 10$ Fock bases. The time-dependent control signal is the detuning $\Delta$ and the amplitude $\alpha_{L}$ of the driven laser within the fixed range $\Delta,\alpha_{L}\in[-5,5]\,\omega_{m}$.
@@
Representative results are as follows. In the target-generating phase, despite the disturbance of the stochastic process from WCM, the training curves for both the reward $\widetilde{R}$ and the logarithmic negativity $\widetilde{E}_{N}$ converge with the episode number, as shown in Figs. 6(a,b), indicating that entanglement has been created and stabilized by the well-trained PPO agent, as shown by Fig. 6(c) with the laser control signal displayed in Fig. 6(d). At the end of the time period, the photon and phonon statistics with respect to the Fock basis are shown in Figs. 6(e,f), where the reduced photon state exhibits an oscillating tail that resembles the displaced squeezed state and the reduced phonon state displays the thermal-like state. Figure 6(g) shows the corresponding target pattern $\langle\hat{n}^{\textnormal{target}}_{p}\rangle(t)$. In the target-utilization phase, the recurrent PPO agent is able to steadily learn to create and stabilize the entanglement, as shown in Fig. 7, where only partial information is extracted from the quantum optomechanical environment. Especially, various entangled states have been created, such as a reduced photon state with the head oscillating on the Fock basis in photon statistics entangled with the thermal-like reduced phonon state, as exemplified in Figs. 7(e,f). Due to the nonlinear and stochastic process in the SME, the entangled states created and controlled are not steady states, rendering infeasible Bayesian control. We thus employ random control as a benchmark, where a uniformly random distribution of actions is taken in a certain range $\Delta,\alpha_{L}\in[-5,5]\omega_{m}$ and the tested values of the measurement rate are $\eta=[0.05,0.1,0.3,0.5,0.7,1]$. Figure 8 shows that, as the measurement rate increases, the random control is unable to harness the entanglement while our well-trained recurrent PPO agent can maintain the entanglement percentage at $50\%$ or higher.
@@
Physical understanding of entanglement engineering through model-free deep RL
@@
Linear interaction regime
For the linear quantum optomechanical system, the maximum entanglement corresponds to a Bell state, of which the expected photon number is $\langle\hat{n}_{p}\rangle=0.5$. Intrinsically, the beam splitter Hamiltonian is capable of generating Bell states [78, 79, 80], a reasonable assumption is that, when the expected photon number reaches the value of 0.5, the maximum entanglement is achieved in a linearly interacting quantum optomechanical system. This assumption provides the base for constructing the reward function $R(t)=-|\langle\hat{n}_{p}\rangle(t)-0.5|$, where the deviation in the expected photon number from 0.5 results in a decreasing reward and therefore implies reduced entanglement. As illustrated in Fig. 9, the RL agent is designed to maximize the accumulated reward value, which is equivalent to stabilizing the expected photon number about the value of 0.5 for as long as possible. The testing results shown in Fig. 3 indicate that the maximum entanglement can indeed be created and stabilized by the RL control.
@@
A central step in RL control is to modulate the laser input based on the measured photon number, which requires the relationship between the laser driving and the photon number. When the frequency of the laser is in the red-detuned regime: $\Delta=\omega_{L}-\omega_{c}=-\omega_{m}$, the quantum state switches between the two modes - the cavity optical and the mechanical oscillator modes, leading to a “swap” Hamiltonian. The coefficient $G$ is proportional to the amplitude of the cavity parameter $\bar{\alpha}_{c}$ that is determined by the laser. In the linear interaction regime, RL control is achieved via two adjustments of the laser based on the measured photon number: (1) the laser frequency is changed into the red-detuned regime and (2) the laser amplitude is perturbed to modulate the driving strength $G$ to control the two modes of switching, which affects the expected photon number. Note that, during this process, there is no energy gain: there is energy loss due to the dissipation of the cavity and oscillator modes into the vacuum bath with the dissipation rate given by $\gamma=0.01\,\kappa$. This relation means that the energy loss due to the oscillator mode occurs more slowly than that with the cavity mode. In essence, the working of the laser is to transfer the energy from the oscillator mode to the cavity mode to stabilize the photon number to a desired value. The underlying dissipation process is not beneficial to the entanglement, as it cannot be modulated by the “swap” term in the Hamiltonian, eliminating any possibility of entanglement enhancement in an optomechanical system in the linear interaction regime. It is worth noting that, in the nonlinear interaction regime, entanglement enhancement and dissipation reduction are possible, as will be described below.
Nonlinear interaction regime
@@
When the interactions between the optical and mechanical modes are nonlinear, the relationship between entanglement and photon number can be sophisticated and is currently unknown. However, model-free deep RL can be used to find the relation numerically. To achieve this, we first assume that there is a solution of the one-to-one correspondence between $E_{N}$ and $\langle\hat{n}_{p}\rangle$ in the time domain. The reward function is constructed according to the target entanglement $E_{N}=\log 2$ to train the RL agent to maximize the accumulated reward. In the testing phase, the time-dependent series of the expected photon number controlled by the well-trained PPO in Fig. 6(g) is regarded as the target time series of the expected photon number for the next target-utilization phase. Note that the “best” photon number is no longer simply 0.5: it is now time-dependent. In the next training phase, the reset RL agent will learn to control the system with the observation $\langle\hat{n}_{p}\rangle(t)$ based on the target’s expected photon number $\langle\hat{n}^{\textnormal{target}}_{p}\rangle(t)$. The performance of the new RL agent in the testing phase, as shown in Fig. 7, validates our initial assumption about the existence of the one-to-one correspondence between $E_{N}$ and $\langle\hat{n}_{p}\rangle$, even though it is time-dependent.
@@
In the nonlinear interaction regime, the physical picture of how the laser leverages the radiation-pressure interaction to create and stabilize the photon number and even the entanglement is not straightforward. However, physical insights can be gained by examining the strong laser limit. When the amplitude of the laser is strong: $\bar{\alpha}_{c}\gg 1$, in the blue-detuned regime with $\Delta=+\omega_{m}$, the laser can modulate its frequency to create exponential growth of the energies of both the cavity and oscillator modes, accompanied by the generation of strong quantum correlation between the two modes. In the red-detuned regime with $\Delta=-\omega_{m}$, a switching process between the two modes occurs, which is the same as that in the linear interaction regime.
@@
The blue- and red-detuned regimes have a competitive relationship with each other in terms of both the photon number and entanglement. In particular, in the blue-detuned regime, photons are excited and the rate of excitation can be larger than that associated with quantum dissipation to the vacuum bath. Furthermore, quantum entanglement is enhanced, overcoming quantum decoherence from the classical environment and even from the SME. However, in the red-detuned regime, no photons are excited and there is only a two-mode energy-transferring process that does not completely suppress the process of quantum dissipation to the vacuum bath, resulting in photon loss and eventually reducing entanglement. Stabilizing the photon number and entanglement requires a balance between the operations in the blue- and red-detuned regimes. In general, the blue-detuned regime is prone to too high photon levels with strong entanglement, which should be balanced by the red-detuned regime operation to reduce the photon number to realize our target entanglement engineering, as shown schematically in Fig. 10. Overall, in the nonlinear interaction regime, laser driving of finite amplitude and frequency modulation can control the photon number and entanglement to a certain extent. An example is shown in Fig. 7(d), where the RL agent finds the optimal action flow with a finite laser amplitude. Note that, the detuning $\Delta$ is modulated mainly in the range $\Delta\in[-2\omega_{m},2\omega_{m}]$, signifying a balance between the blue- and red-detuned operations.
Weak continuous measurement
@@
In an open quantum system, under WCM and quantum dissipation into the vacuum bath as well, a Wiener process occurs in the observable. More specifically, the Wiener process arises from the Gaussian-weighted projection over the eigenstates, which weakly extracts the partial information from the quantum system and induces stochastic disturbances in both the dynamical equation and observation. Such disturbances can avoid a complete quantum state collapse and provide the capability to extract the quantum information continuously in the time domain. However, the nonlinear stochastic process occurs in both quantum dynamical trajectories and the measurement photocurrent, making it challenging to control the quantum system through WCM continuously. (Backgrounds about WCM, deep RL, and quantum control are presented in Supplementary Note 1.)
@@
For stochastic noise in the WCM photocurrent, the present cutting-edge technology enables the RL agent to extract quantum information through a process resembling noise filtering. Specifically, the observation in the reward function is the WCM photocurrent. We can employ $n_{\textnormal{traj}}$ quantum ensembles to reduce the variance and use a Gaussian filter for data pre-processing. The RL agent is trained to maximize the accumulated reward, which serves to average the stochastic term in the measurement photocurrent over time. These noise-filtering processes help extract information about the expected photon number and thus the target quantum entanglement. For the nonlinear quantum stochastic process with quantum dissipation, the RL agent successfully trains the laser to leverage interactions between the optical and mechanical modes, linear or nonlinear, to mitigate quantum decoherence and dissipation to some extent, as exemplified in Figs. 3 and 7.
Discussion